Alice needs a return of 12% on her investment. She has $14,000 in bonds that yield 8% and $16,000 to invest. What yield should she get on the $16,000?
Answer by mananth(16946)

Total investment = > 14000+16000=30,000
return = 12%
=3,600
..
14,000 @ 8% = 1120
total return - return on bonds = 3600-1120=2480
..
She has 16000 to invest
return required = 2480
I= p*r*t
1480 = 16000*r*1
2480/16000=r
r= 0.155
= 15.5%
she has to invest at 15.5%